# Seeing through Uncertainty: Robust Task-Oriented Optimization in Visual Navigation

### v1

@@ 1 Introduction
Figure 1: Overview of the work. (Left) Traditional network-based agents exhibit limited generalization capabilities in visual navigation. (Right) The foundational concept of a task-based optimization framework and its inherent challenges when directly applied to visual navigation. NeuRO framework addresses this mismatch through a conformal visual-processing method and a robust optimization formulation for decision-making.
@@ 4.1 Comparison with SoTA
Figure 3: Learning curves for NeuRO and baseline during training for different tasks. $m=2$.
@@ Appendix G Broader Application: Power Grid Scheduling as a Case Study
The task loss and prediction loss for NeuRO and ETO were recorded and are shown in Table˜6, and the visualization of prediction loss can be found in Fig.˜6 (the grey region is the visualization of the box uncertainty from the estimation). Table˜6 indicates that NeuRO’s prediction accuracy is slightly lower than the traditional approach, but its task performance improves by 77%. Therefore, we believe NeuRO’s learning ability inherently involves a trade-off. However, in real-world applications, only accurate price predictions hold practical significance, making NeuRO more suitable for tasks where precise recognition is not required. For example, in navigation, we care less about whether the robot forms human-like cognition and more about how it interacts with the environment to complete tasks.
Appendix H Additional Experimental Results
Agent Performance During Training
We analyzed the training progression of NeuRO by comparing its learning curves against several baselines: (i) NeuRO without its task-specific optimization component (see main paper Ablation Study Section for more about Task-Specific Optimization), (ii) a previous SoTA method Lyon, and (iii) our foundational baseline, OracleEgoMap. (Fig.˜7) Across various navigation tasks,NeuROconsistently exhibited faster convergence rates and achieved superior final task performance compared to these alternatives. These training dynamics further substantiate the benefits of integrating explicit optimization within the learning agent, leading to more efficient and effective policy acquisition.
Figure 7: Evaluation metrics of agents during training with two seeds.
Generalization Across Task Variations
To assess the generalization capabilities of our agents, we conducted experiments where models trained on a specific instance of the S-MONtask, denoted S-MONm=i (where $m$ represents the number of navigation goals, and $i$ is the specific number of sub-goals used for training), were evaluated on different instances SMONm=j. The performance, typically measured by success rate, is reported in Table 7.
OracleEgoMap Lyon (SoTA) NeuRO
[…]
1 83 47 28 86 61 50 90 68 57
2 79 64 32 82 76 54 88 80 60
3 77 63 37 81 75 57 88 79 62
1 0 -17 -9 0 -15 -7 0 -12 -5
2 -4 0 -5 -4 0 -3 -2 0 -2
3 -6 -1 0 -5 -1 0 -2 -1 0
H.1 Impact of Target Object Geometry

### v3

@@ 1 Introduction
Figure 1: Overview of the work. (Left) The foundational concept of a task-based optimization framework and its inherent challenges when directly applied to visual navigation. (Right) NeuRO framework addresses this mismatch through a conformal visual-processing method and a robust optimization formulation for decision-making.
@@ 4.2 Quantitative and Qualitative Study
Figure 3: Learning curves for NeuRO and baseline during training for different tasks.
@@ 4.3 Ablation Studies
Generalization Across Task Variations. To assess the generalization capabilities of our agents, we conducted experiments where models trained on a specific instance of the S-MONtask, denoted S-MONm=i (where $m$ represents the number of navigation goals, and $i$ is the specific number of sub-goals used for training), were evaluated on different instances SMONm=j. The performance, typically measured by success rate, is reported in Table 5.
OracleEgoMap Lyon (SoTA) NeuRO
[…]
1 83 47 28 86 61 50 90 68 57
2 79 64 32 82 76 54 88 80 60
3 77 63 37 81 75 57 88 79 62
1 0 -17 -9 0 -15 -7 0 -12 -5
2 -4 0 -5 -4 0 -3 -2 0 -2
3 -6 -1 0 -5 -1 0 -2 -1 0
@@ 5 Conclusion
Acknowledgment
This paper was supported by the Natural Science Foundation of China under Grant 62303307, 62225309, U24A20278, 62361166632, U21A20480, and Sponsored by the Oceanic Interdisciplinary Program of Shanghai Jiao Tong University, and in part by National Key Laboratory of Human Machine Hybrid Augmented Intelligence, Xi’an Jiaotong University (No. HMHAI-202408).
@@ Appendix G Broader Application: Power Grid Scheduling as a Case Study
The task loss and prediction loss for NeuRO and ETO were recorded and are shown in Table˜7, and the visualization of prediction loss can be found in Fig.˜6 (the grey region is the visualization of the box uncertainty from the estimation). Table˜7 indicates that NeuRO’s prediction accuracy is slightly lower than the traditional approach, but its task performance improves by 77%. Therefore, we believe NeuRO’s learning ability inherently involves a trade-off. However, in real-world applications, only accurate price predictions hold practical significance, making NeuRO more suitable for tasks where precise recognition is not required. For example, in navigation, we care less about whether the robot forms human-like cognition and more about how it interacts with the environment to complete tasks.
Appendix H Impact of Target Object Geometry